Suppose you start saving $12,000 per year (starting next year) for the 35 years you are working. How much can you withdraw each year for the 40 years you are retired? Assume that you earn an 8.5% return for the years you are saving and a 3.5% return for the years you are retired.
Answer #1

Answer:

Given

Annual Saving S=$12000

number of payments N=35

interest during saving R=8.5%

Number of Withdrawal n=40

interest during withdrawal r=3.5%

Let A be amount can be withdrawn during retirement

So Sum of Future Value of Saving= Sum of Present value of Withdrawals

S*((1+R)^N-1)/R= A*(1-(1+r)^-n)/r

12000*((1+8.5%)^35-1)/8.5%=A*(1-(1+3.5%)^-40)/3.5%

21.36A=2312420.10

A=$108284.35